OpenStudy (anonymous):

find the discriminant for 2x^ - x -6 = 0

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

\[y=2x^2−1x−6\] \[discriminant=b^2−4(a)(c)\]

OpenStudy (kidrah69):

\[(−1)2−4(2)(−6)=?\]

OpenStudy (anonymous):

where did you get the negative one ?

OpenStudy (kidrah69):

-x=-1

OpenStudy (anonymous):

so do i just solve it from there?

OpenStudy (kidrah69):

yep :)

OpenStudy (anonymous):

so i got -47 , and there are none for solutions am i right?

OpenStudy (kidrah69):

hnnn thats not right.

OpenStudy (anonymous):

oh.. then what is it

OpenStudy (kidrah69):

Im not giving you the answer Solve this (−1)2−4(2)(−6)

OpenStudy (anonymous):

-50

OpenStudy (kidrah69):

Still wrong.

OpenStudy (anonymous):

what am i doing wrong am i subtracting wrong ?

OpenStudy (kidrah69):

ok (-1)^2=?

OpenStudy (anonymous):

= 2 ? haha im confused ;c

OpenStudy (kidrah69):

it would be 1 :P

OpenStudy (kidrah69):

1*2=?

OpenStudy (kidrah69):

y=2x2−1x−6 discriminant=b2−4(a)(c) =(−1)2−4(2)(−6)=49 ( 2 real solutions )

OpenStudy (kidrah69):

Nevermind im confusing u more :) here look at that

OpenStudy (anonymous):

oh. >.< oh my goodness thankyou for ending the confusion